OpenStudy (horsegal244):

What is the sum of 2.5 × 104 and 3.4 × 104?

OpenStudy (horsegal244):

0.59 × 10^4 0.59 × 10^8 5.9 × 10^4 5.9 × 10^8

OpenStudy (alekos):

(a x b) + (c x b) = (a + c) x b

OpenStudy (horsegal244):

@alekos I dont get that...

OpenStudy (horsegal244):

@LynFran

OpenStudy (alekos):

a=2.5 c=3.4 and b =10^4

OpenStudy (horsegal244):

ok

OpenStudy (alekos):

what answer did you get?